# DU-Shapley: A Shapley Value Proxy for Efficient Dataset Valuation

### v2

@@ 1 Introduction
Additional Related Works. Cooperative game theory has been applied to solve multi-agents ML problems beyond data and dataset valuation [6, 18, 27, 44]. In particular, the Shapley value has been used to solve several problems including variable selection [5], feature importance [7, 25, 26], or model interpretation [4]. In these problems, similarly to the data and dataset valuation problems, the computational intractability issue of the Shapley value is usually addressed via MC [2, 29, 30].
@@ 2.1.3 Theoretical use-case 3: Heterogeneous Linear Regression - Local Differential Privacy
where the expectation is taken over the distribution $p_{\mathrm{test}}$ of a hold-out testing datum $x\in\mathbb{R}^{d}$, the sampling distributions $\mathrm{N}(0,\sigma_{i}^{2}\mathrm{I}_{d}),\forall i\in\mathcal{S}$, and the linear regression error distributions $\mathrm{N}(0,\varepsilon_{i}^{2}),\forall i\in\mathcal{S},j\in[n_{i}]$, and $\hat{\theta}_{\mathcal{S}}$ stands for the generalized least square estimator defined by $\hat{\theta}_{\mathcal{S}}=(X_{\mathcal{S}}^{\top}\Sigma_{\mathcal{S}}^{-1}X_{\mathcal{S}})^{-1}X_{\mathcal{S}}^{\top}\Sigma_{\mathcal{S}}^{-1}Y_{\mathcal{S}},\text{ where }\Sigma_{\mathcal{S}}=\mathrm{diag}((\varepsilon^{2}_{i})_{i\in\mathcal{S}})\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}.$ The notations $X_{\mathcal{S}}$ and $Y_{\mathcal{S}}$ refer to the concatenation of $\{X_{i}\}_{i\in\mathcal{S}}$ and $\{Y_{i}\}_{i\in\mathcal{S}}$, respectively, and $X_{i}\in\mathbb{R}^{n_{i}\times d}$ is defined by $X_{i}=([x_{i}^{(1)}]^{\top},\ldots,[x_{i}^{(n_{i})}]^{\top})^{\top}$ while $Y_{i}\in\mathbb{R}^{n_{i}}$ is defined by $Y_{i}=(y_{i}^{(1)},\ldots,y_{i}^{(n_{i})})^{\top}$.
@@
Section 2.1.3 shows that, in this use-case, the utility function can be written as a function $w(q(\mathcal{S}))$ of a scalar quantity $q(\mathcal{S})$ that captures the datasets heterogeneity. Notice that in this use-case, if we add the homogeneity assumption that $\sigma_{i}/\varepsilon_{i}=\sigma/\varepsilon,\forall i\in\mathcal{I}$, then the term $q(\mathcal{S})$ becomes $\sum_{i\in\mathcal{S}}n_{i}$ and, as a consequence, we get
@@ 3.1 Insights behind DU-Shapley
$q(\mathcal{S}):=\biggl{\lfloor}\frac{\bigl{(}\sum_{i\in\mathcal{S}}\gamma_{i}n_{i}\bigr{)}^{2}}{\sum_{i\in\mathcal{S}}\gamma_{i}^{2}n_{i}}\biggr{\rfloor},\text{ where }\forall i\in\mathcal{I},\gamma_{i}=\left\{\begin{array}[]{cl}1&\text{for the second use-case},\\ \nicefrac{{\sigma_{i}}}{{\varepsilon_{i}}}&\text{for the third use-case},\end{array}\right.$ (14)
@@ 3.2 Discrete Uniform Shapley value
$\psi_{i}(u):=\frac{1}{I}\sum\nolimits_{k=0}^{I-1}u(\mathrm{D}^{(k)}\cup\mathrm{D}^{(k)}_{i})-u(\mathrm{D}^{(k)}),$ (17)
@@
The sampling procedure of DU-Shapley can be significantly improved whenever $u(\mathcal{S})=w(q(\mathcal{S}))$, with $q(\mathcal{S})$ some scalar quantify of interest, as in our use-cases. Indeed, instead of pooling the datasets together and sampling datapoints, it is enough to consider $I$ real values from well chosen intervals. We rewrite DU-Shapley as
$\psi_{i}(w)=\frac{1}{I}\sum\nolimits_{k=1}^{I-1}w(k\bar{q})-w(k\bar{q}_{-i})+\Delta_{0}^{i},$ (18)
where $\bar{q}=q(\mathcal{I)}/(I-1)$, $\bar{q}_{-i}=q(\mathcal{I}\setminus\{i\})/(I-1)$, $\Delta_{0}^{i}=u(\{i\})-u(\varnothing)$ is the offset term, and $q(\cdot)$ is defined in Equation (14). We obtain the following result from Theorem 1.
@@
While our theoretical results are based on the simplified DU-Shapley (18) for the cases where $u(\mathcal{S})=w(q(\mathcal{S}))$, we will see through numerical experiments that Section 3.2 gives good results in more general cases.
@@ 4 Numerical Experiments
We illustrate the benefits of DU-Shapley by measuring numerically two properties on real data: (1) how well DU-Shapley approximates the Shapley value and (2) how well DU-Shapley performs in classical dataset valuation tasks. Appendices A and B.2 complement the results by a computational complexity comparison between our method and MC and experiments on synthetic data, respectively. The experiments strongly suggest that DU-Shapley performs well in all tasks.
@@ 4.1 Approximating the Shapley value in Real-World Data
Dataset $I$ DU-Shapley MC-Shapley MC-anti-Shapley
adult 10 $\boldsymbol{2.10^{-3}}$ $1.10^{-2}$ $8.10^{-3}$
adult 20 $\boldsymbol{6.10^{-4}}$ $4.10^{-3}$ $2.10^{-3}$
breast-cancer 10 $\boldsymbol{3.10^{-3}}$ $3.10^{-2}$ $1.10^{-2}$
breast-cancer 20 $\boldsymbol{1.10^{-4}}$ $1.10^{-3}$ $8.10^{-4}$
bank 10 $\boldsymbol{5.10^{-2}}$ $9.10^{-2}$ $8.10^{-2}$
bank 20 $\boldsymbol{4.10^{-3}}$ $6.10^{-2}$ $4.10^{-2}$
cal-housing 10 $\boldsymbol{1.10^{-2}}$ $5.10^{-2}$ $3.10^{-2}$
cal-housing 20 $\boldsymbol{3.10^{-3}}$ $2.10^{-2}$ $1.10^{-2}$
make-regression 10 $\boldsymbol{9.10^{-2}}$ $4.10^{-1}$ $4.10^{-1}$
make-regression 20 $\boldsymbol{2.10^{-2}}$ $3.10^{-1}$ $2.10^{-1}$
year 10 $\boldsymbol{1.10^{-3}}$ $5.10^{-3}$ $5.10^{-3}$
year 20 $\boldsymbol{7.10^{-4}}$ $1.10^{-3}$ $1.10^{-3}$
@@ 4.2 Applying DU-Shapley to dataset valuation problems
We considered non-tabular datasets used in OpenDataVal, namely bbc-embedding, IMDB-embedding, both text datasets, and CIFAR10-embedding, an image dataset. Feature embedding have been generated using pretrained DistilBERT and ResNet50 models, respectively. In addition we have adapted three baselines from data valuation to our setting: Leave-One-Out (LOO), DataShapley, and KNN-Shapley. Section B.4 gives the implementations details. For these datasets associated to classification problems, we used a multi-layer perceptron classifier as prediction model.
We have considered three dataset valuation problems, none of them needing the real Shapley values, which allows us to increase the number of players w.r.t. the experiments in Section 4.1. We investigated noisy label detection (NLD), dataset removal (DR), and dataset addition (DA) [17]. For NLD, we used as a metric the F1-score (the larger the better). For DR, we used the testing accuracy (the larger the better). For DA, we used the testing accuracy (the lesser the better).
@@
Table 2: Comparison between DU-Shapley and competitors for real-world datasets considered in OpenDataVal in Noisy label detection, Dataset Removal and Dataset Addition.
Dataset Problem Random LOO Data Shapley KNN-Shapley DU-Shapley
CIFAR 10 NLD 5% 0.11 0.13 0.13 0.14 0.14
CIFAR 10 NLD 15% 0.19 0.18 0.25 0.28 0.30
CIFAR 10 DR 5% 0.61 0.62 0.61 0.60 0.61
CIFAR 10 DR 15% 0.60 0.60 0.59 0.57 0.55
CIFAR 10 DA 5% 0.25 0.15 0.12 0.12 0.11
CIFAR 10 DA 15% 0.41 0.32 0.18 0.15 0.14
BBC NLD 5% 0.11 0.11 0.12 0.19 0.18
BBC NLD 15% 0.19 0.17 0.20 0.29 0.34
BBC DR 5% 0.90 0.90 0.89 0.88 0.89
BBC DR 15% 0.88 0.88 0.87 0.86 0.85
BBC DA 5% 0.68 0.61 0.08 0.13 0.07
BBC DA 15% 0.81 0.77 0.12 0.12 0.11
IMDB NLD 5% 0.10 0.11 0.17 0.18 0.18
IMDB NLD 15% 0.16 0.18 0.28 0.29 0.32
IMDB DR 5% 0.77 0.77 0.75 0.76 0.76
IMDB DR 15% 0.75 0.74 0.69 0.68 0.66
IMDB DA 5% 0.62 0.53 0.36 0.41 0.33
IMDB DA 15% 0.68 0.59 0.33 0.37 0.34
@@ 5 Conclusion
We model the dataset valuation problem as a cooperative game and design a Shapley value approximation, named DU-Shapley, that exploits the underlying structure of the utility function and exponentially reduces the number of functions valuations required for the computation. In three different uses-cases, DU-Shapley is proved to almost surely converge to the real Shapley value as the number of players grows. Moreover, we find the rate of convergence, which depends only on natural parameters of dataset valuation. Numerical experiments showcase that DU-Shapley performs well in approximating the Shapley value and performing dataset valuation tasks, even when the assumptions needed for the theoretical guarantees do not hold.
@@
Appendix A Complexity comparison of DU-Shapley and Monte Carlo
It is crucial to notice that computing the DU-Shapley of a player $i\in\mathcal{I}$ requires $I$ function evaluations. Regarding Monte Carlo, Hoeffding’s result [14] implies that a lower bound on the number of permutations $T$, such that $\mathbb{P}(\|\varphi_{i}(u)-\hat{\varphi}_{i}(u)\|_{2}\leq\varepsilon)\geq 1-\delta$, is given by $T_{\text{perm}}(\varepsilon,\delta)=(2r_{u}^{2}I/\varepsilon^{2})\log(2I/\delta)$, where $\varepsilon,\delta\in(0,1)$ and $r_{u}=\max_{\mathcal{S}_{1},\mathcal{S}_{2}\subseteq\mathcal{I}}\{|u(\mathcal{S}_{1})-u(\mathcal{S}_{2})|\}$ is the range of the utility function $u$. Assuming the existence of a function $w:\mathbb{R}_{+}\to\mathbb{R}$ as in the uses-cases and fixing the sampling budget to $I$, i.e., equal to the number of terms needed to compute DU-Shapley, the Monte Carlo approximation error becomes
$\varepsilon(I)=2w(n_{\mathcal{I}})^{2}\log\biggl{(}\frac{2I}{\delta}\biggr{)},$ (22)
where we have supposed that $0=w(n_{\varnothing})\leq w(n_{\mathcal{S}})\leq w(n_{\mathcal{I}})$ for any $\mathcal{S}\subseteq\mathcal{I}$, so that $r_{w}=w(n_{\mathcal{I}})$. Notice that $\varepsilon(I)$ is increasing with $I$, while the bias of DU-Shapley in Theorem 2 decreases as the number of players grows. Figure 2 compares (22) with DU-Shapley’s bias (19) as the number of players increases, in a mixed between our second and third use-cases [^2] The value function used in Figure 2 is motivated by the homogeneous version of our third use-case and taken this way to obtain a positive and increasing function. Under no circumstances the conclusions of Figure 2 are influenced by this choice.. By considering the same computational budget for both approaches, we can observe that the Monte Carlo approximation of the Shapley value is associated with a larger approximation error than DU-Shapley, when the number of players becomes sufficiently large. This result confirms that DU-Shapley is indeed relevant for dataset valuation in this regime.
Figure 2: Monte Carlo’s expected error for limited sampling budget ($T=I$) versus DU-Shapley’s expected bias for a value function $w(n)=1-\frac{10^{k(\mathcal{I})}}{10^{k(\mathcal{I})}+n}$ where $k(\mathcal{I})=\lfloor\log(\sum_{i\in\mathcal{I}}n_{i})\rfloor-1$. For each value of $I$, we drew $100$ times the data points of each player from $\mathrm{U}([n_{\mathrm{max}}])$, with (left) $n_{\mathrm{max}}=10^{2}$, (middle) $n_{\mathrm{max}}=10^{3}$, and (right) $n_{\mathrm{max}}=10^{4}$.
Appendix B Complementary Numerical results

### v3

@@ 1 Introduction
Additional Related Work. Cooperative game theory has been applied to solve multi-agents ML problems beyond data and dataset valuation [6, 18, 29, 47]. In particular, the Shapley value has been used to solve several problems including variable selection [5], feature importance [7, 27, 28], or model interpretation [4]. In these problems, similarly to the data and dataset valuation problems, the computational intractability issue of the Shapley value is usually addressed via MC [2, 31, 32].
@@ 2.1.3 Theoretical use-case 3: Heterogeneous Linear Regression - Local Differential Privacy
where the expectation is taken over the distribution $p_{\mathrm{test}}$ of a hold-out testing datum $x\in\mathbb{R}^{d}$, the sampling distributions $\mathrm{N}(0,\sigma_{i}^{2}\mathrm{I}_{d})$ for all $i\in\mathcal{S}$, and the linear regression error distributions $\mathrm{N}(0,\varepsilon_{i}^{2}),\forall i\in\mathcal{S},j\in[n_{i}]$, and $\hat{\theta}_{\mathcal{S}}$ stands for the generalized least square estimator defined by $\hat{\theta}_{\mathcal{S}}=(X_{\mathcal{S}}^{\top}\Sigma_{\mathcal{S}}^{-1}X_{\mathcal{S}})^{-1}X_{\mathcal{S}}^{\top}\Sigma_{\mathcal{S}}^{-1}Y_{\mathcal{S}},\text{ where }\Sigma_{\mathcal{S}}=\mathrm{diag}((\varepsilon^{2}_{i})_{i\in\mathcal{S}})\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}.$ The notations $X_{\mathcal{S}}$ and $Y_{\mathcal{S}}$ refer to the concatenation of $\{X_{i}\}_{i\in\mathcal{S}}$ and $\{Y_{i}\}_{i\in\mathcal{S}}$, respectively, and $X_{i}\in\mathbb{R}^{n_{i}\times d}$ is defined by $X_{i}=([x_{i}^{(1)}]^{\top},\ldots,[x_{i}^{(n_{i})}]^{\top})^{\top}$ while $Y_{i}\in\mathbb{R}^{n_{i}}$ is defined by $Y_{i}=(y_{i}^{(1)},\ldots,y_{i}^{(n_{i})})^{\top}$.
@@
Section 2.1.3 shows that, in this use-case, the utility function can be written as a function $w(q(\mathcal{S}))$ of a scalar quantity $q(\mathcal{S})$ that captures the datasets heterogeneity. Notice that in this use-case, if we add the homogeneity assumption that $\sigma_{i}/\varepsilon_{i}=\sigma/\varepsilon$, for all $i\in\mathcal{I}$, then the term $q(\mathcal{S})$ becomes $\sum_{i\in\mathcal{S}}n_{i}$ and, as a consequence, we get
@@ 3.1 Insights behind DU-Shapley
$q(\mathcal{S}):=\biggl{\lfloor}\frac{\bigl{(}\sum_{i\in\mathcal{S}}\gamma_{i}n_{i}\bigr{)}^{2}}{\sum_{i\in\mathcal{S}}\gamma_{i}^{2}n_{i}}\biggr{\rfloor},\text{ where, for any }i\in\mathcal{I},\gamma_{i}=\left\{\begin{array}[]{cl}1&\text{for the second use-case},\\ \nicefrac{{\sigma_{i}}}{{\varepsilon_{i}}}&\text{for the third use-case},\end{array}\right.$ (14)
@@ 3.2 Discrete Uniform Shapley value
$\psi_{i}(u):=\frac{1}{I}\sum\nolimits_{k=0}^{I-1}u(\mathrm{D}^{(k)}\cup\mathrm{D}_{i})-u(\mathrm{D}^{(k)}),$ (17)
@@
By definition, DU-Shapley is a random variable which depends on the sampled data points. However, whenever $u(\mathcal{S})=w(q(\mathcal{S}))$, with $q(\mathcal{S})$ some scalar quantify of interest, as in our use-cases, we can get rid of the stochastic nature of DU-Shapley by considering $I$ real values from well chosen intervals. In particular, in our uses-cases, DU-Shapley boils down to:
$\psi_{i}(w)=\frac{1}{I}\sum\nolimits_{k=0}^{I-1}w(\bar{q}_{i}^{k})-w(\bar{q}_{-i}^{k}),$ (18)
where
$\bar{q}_{i}^{k}:=\biggl{\lfloor}\frac{(\gamma_{i}n_{i}+\frac{k}{I-1}\sum_{j\in\mathcal{I}\setminus\{i\}}\gamma_{j}n_{j})^{2}}{\gamma_{i}^{2}n_{i}+\frac{k}{I-1}\sum_{j\in\mathcal{I}\setminus\{i\}}\gamma_{j}^{2}n_{j}}\biggr{\rfloor}\text{ and }\bar{q}_{-i}^{k}:=\biggl{\lfloor}\frac{k}{I-1}\cdot\frac{(\sum_{j\in\mathcal{I}\setminus\{i\}}\gamma_{j}n_{j})^{2}}{\sum_{j\in\mathcal{I}\setminus\{i\}}\gamma_{j}^{2}n_{j}}\biggr{\rfloor}.$
We remark the notation abuse as we should write $\psi(w\circ q)$. For simplicity, we omit the composition and only write $\psi(w)$. Equation 18 coincides exactly with Section 3.2 in the first two use-cases, i.e., when $\gamma_{j}=\gamma$ for all $j\in\mathcal{I}$. Indeed, as the random datasets $\mathrm{D}^{(k)}$ have a fixed size and the value function only looks at the number of data points within the coalition, we obtain,
$\psi_{i}(u)$ $=\frac{1}{I}\sum_{k=0}^{I-1}u(\mathrm{D}^{(k)}\cup\mathrm{D}_{i})-u(\mathrm{D}^{(k)})=\frac{1}{I}\sum_{k=0}^{I-1}w(|\mathrm{D}^{(k)}\cup\mathrm{D}_{i}|)-w(|\mathrm{D}^{(k)}|)$ (19)
$=\frac{1}{I}\sum_{k=0}^{I-1}w(k\mu_{-i}+n_{i})-w(k\mu_{-i})=\psi_{i}(w).$ (20)
For the third use-case, Equation 18 is an approximation that comes from assuming that, for any $j\in\mathcal{I}\setminus\{i\}$, $|\mathrm{D}_{j}\cap\mathrm{D}^{(k)}|=k\cdot\frac{n_{j}}{I-1}$, which holds with high probability for large values of $I$, since
$q(\mathrm{D}^{(k)}\cup\mathrm{D}_{i})=\biggl{\lfloor}\frac{\bigl{(}\gamma_{i}n_{i}+\sum_{j\in\mathcal{I}\setminus\{i\}}\gamma_{j}\cdot|\mathrm{D}_{j}\cap\mathrm{D}^{(k)}|\bigr{)}^{2}}{\gamma_{i}^{2}n_{i}+\sum_{j\in\mathcal{I}\setminus\{i\}}\gamma_{j}^{2}\cdot|\mathrm{D}_{j}\cap\mathrm{D}^{(k)}|}\biggr{\rfloor}.$ (21)
Theorem 1 implies the following result.
@@
While our theoretical results are based on Equation (18) for the cases where $u(\mathcal{S})=w(q(\mathcal{S}))$, we will see through numerical experiments that Section 3.2 gives good results in more general cases.
@@ 4 Numerical Experiments
We illustrate the benefits of DU-Shapley by measuring numerically three properties: (1) how well DU-Shapley approximates the Shapley value in real data, (2) how many (theoretical) iterations need other methods to achieve the same accuracy level than DU-Shapley, and (3) how well DU-Shapley performs in classical dataset valuation tasks with real data. Section A.1 complements the results by a complexity comparison between our method and SVARM [22] and Section A.2 by experiments on synthetic data. The experiments strongly suggest that DU-Shapley performs well in all tasks.
@@ 4.1 Approximating the Shapley Value in Real-World Data
Dataset adult breast-cancer bank cal-housing
Players 10 20 10 20 10 20 10 20
DU-Shapley $\mathbf{2.10^{-3}}$ $\mathbf{6.10^{-4}}$ $\mathbf{3.10^{-3}}$ $\mathbf{1.10^{-4}}$ $\mathbf{5.10^{-2}}$ $\mathbf{4.10^{-3}}$ $\mathbf{1.10^{-2}}$ $\mathbf{3.10^{-3}}$
MC-Shapley $1.10^{-2}$ $4.10^{-3}$ $3.10^{-2}$ $1.10^{-3}$ $9.10^{-2}$ $6.10^{-2}$ $5.10^{-2}$ $2.10^{-2}$
MC-anti-Shapley $8.10^{-3}$ $2.10^{-3}$ $1.10^{-2}$ $8.10^{-4}$ $8.10^{-2}$ $4.10^{-2}$ $3.10^{-2}$ $1.10^{-2}$
Dataset make-regression year
Players 10 20 10 20
DU-Shapley $\mathbf{9.10^{-2}}$ $\mathbf{2.10^{-2}}$ $\mathbf{1.10^{-3}}$ $\mathbf{7.10^{-4}}$
MC-Shapley $4.10^{-1}$ $3.10^{-1}$ $5.10^{-3}$ $1.10^{-3}$
MC-anti-Shapley $4.10^{-1}$ $2.10^{-1}$ $5.10^{-3}$ $1.10^{-3}$
4.2 Complexity of Computing the Shapley Values of all Players
We have looked at the number of iterations that DataShapley and the Improved Group Testing-Based method [46] (IGTB) require to achieve DU-Shapley’s accumulated bias, formally given by
$\mathrm{DUbias}(I):=\frac{\kappa}{I-1}\biggl{(}\sum\nolimits_{i\in\mathcal{I}}\frac{(9\sigma_{-i}^{2}(1+\log(I-1))+\zeta_{-i})^{2}}{(\mu_{-i})^{4}}\biggr{)}^{1/2}.$ (25)
To do so, we have replaced $\varepsilon=\mathrm{DUbias}(I)$, respectively, in the formula in Section 4.1 in [16] and Equation 5 in [46], with a value function motivated from our third use-case under the homogeneity assumption $\sigma_{i}/\varepsilon_{i}=\sigma/\varepsilon$ for all $i\in\mathcal{I}$. The results are illustrated in Figure 2. Remark DU-Shapley requires $I^{2}$ iterations to compute all Shapley values. We observe that in all tested instances, both methods require a higher number of iterations to achieve the same error than DU-Shapley.
Figure 2: Iterations required by DataShapley and the Improved Group Testing-Based method to achieve DU-Shapley’s accumulated bias with function $w(n_{S})=1-\frac{10^{k(\mathcal{I})}}{10^{k(\mathcal{I})}+n_{S}}$, where $n_{S}$ is the number of data points of the coalition $S\subseteq\mathcal{I}$, and $k(\mathcal{I}):=\lfloor\log(\sum_{i\in\mathcal{I}}n_{i})\rfloor-1$ is a normalization factor. (top) $\delta=0.01$, (bottom) $\delta=0.1$, (left) $n_{\mathrm{max}}=10$, (middle) $n_{\mathrm{max}}=50$, (right) $n_{\mathrm{max}}=100$.
@@ 4.3 Applying DU-Shapley to dataset valuation problems
We considered non-tabular datasets used in [17], namely bbc-embedding, IMDB-embedding, both text datasets, and CIFAR10-embedding, an image dataset. Feature embedding have been generated using pretrained DistilBERT and ResNet50 models, respectively. In addition we have adapted three baselines from data valuation to our setting: Leave-One-Out (LOO), DataShapley, and KNN-Shapley. Section B.2 gives the implementations details. For these datasets associated to classification problems, we used a multi-layer perceptron classifier as prediction model.
We have considered three dataset valuation problems, none of them needing the real Shapley values, which allows us to increase the number of players w.r.t. the experiments in Section 4.1. We investigated noisy label detection (NLD), dataset removal (DR), and dataset addition (DA) [17]. For NLD, we used as a metric the F1-score (the larger the better). For DR, we used the testing accuracy (the lesser the better). For DA, we used the testing accuracy (the lesser the better).
@@
Table 2: Comparison between DU-Shapley and competitors for real-world datasets considered in [17] in Noisy label detection, Dataset Removal and Dataset Addition.
Dataset CIFAR 10 BBC
Problem NLD DR DA NLD DR DA
5% 15% 5% 15% 5% 15% 5% 15% 5% 15% 5% 15%
Random 0.11 0.19 0.61 0.60 0.25 0.41 0.11 0.19 0.90 0.88 0.68 0.81
LOO 0.13 0.18 0.62 0.60 0.15 0.32 0.11 0.17 0.90 0.88 0.61 0.77
DataShapley 0.13 0.25 0.61 0.59 0.12 0.18 0.12 0.20 0.89 0.87 0.08 0.12
KNN-Shapley 0.14 0.28 0.60 0.57 0.12 0.15 0.19 0.29 0.88 0.86 0.13 0.12
DU-Shapley 0.14 0.30 0.61 0.55 0.11 0.14 0.18 0.34 0.89 0.85 0.07 0.11
Dataset IMBD
Problem NLD DR DA
5% 15% 5% 15% 5% 15%
Random 0.10 0.16 0.77 0.75 0.62 0.68
LOO 0.11 0.18 0.77 0.74 0.53 0.59
DataShapley 0.17 0.28 0.75 0.69 0.36 0.33
KNN-Shapley 0.18 0.29 0.76 0.68 0.41 0.37
DU-Shapley 0.18 0.32 0.76 0.66 0.33 0.34
@@ 5 Conclusion
We model the dataset valuation problem as a cooperative game and design a Shapley value approximation, named DU-Shapley, that exploits the underlying structure of the utility function and exponentially reduces the number of functions valuations required for the computation. In three different uses-cases, DU-Shapley is proved to almost surely converge to the real Shapley value as the number of players grows. Moreover, we find the rate of convergence, which depends only on natural parameters of dataset valuation. Numerical experiments showcase that DU-Shapley performs well in approximating the Shapley value and performing dataset valuation tasks, even when the assumptions needed for the theoretical guarantees do not hold, and it has a good complexity when computing the Shapley values of all players.
@@
Acknowledgments
This research was supported in part by the French National Research Agency (ANR) in the framework of the PEPR IA FOUNDRY project (ANR-23-PEIA-0003) and through the grant DOOM ANR-23-CE23-0002. It was also funded by the European Union (ERC, Ocean, 101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Complementary Numerical results